# Revisiting Shooting Point Monte Carlo Methods for Transition Path Sampling

### v1

@@ Abstract
Rare event sampling algorithms are essential for understanding processes that occur infrequently on the molecular scale, yet they are important for the long-time dynamics of complex molecular systems. One of these algorithms, transition path sampling, has become a standard technique to study such rare processes since no prior knowledge on the transition region is required. Most TPS methods generate new trajectories from old trajectories by selecting a point along the old trajectory, modifying its momentum in some way, and then “shooting” a new trajectory by integrating forward and backward in time. In some procedures, the shooting point is selected independently for each trial move, but in others, the shooting point evolves from one path to the next so that successive shooting points are related to each other. We provide an extended detailed balance criterion for shooting methods. We affirm detailed balance for most TPS methods, but the new criteria reveals the need for amended acceptance criteria in the flexible length aimless shooting and spring shooting methods.
@@ I Introduction
Transition path sampling (TPS) is a rare events method that yields unbiased insights into the mechanism and kinetics in complex molecular systems, including processes such as nucleation Menzl et al. (2016), biomolecular reorganization ichi Okazaki et al. (2019); Juraszek and Bolhuis (2006) and chemical reactions Basner and Schwartz (2005); Geissler et al. (2001). Unlike many enhanced sampling techniques, TPS does not require detailed a priori knowledge about the process of interest, apart from a definition of the relevant (meta)stable states, i.e., knowledge of the start and end point of the process Bolhuis et al. (2002). The aim of TPS is sampling from a distribution of unbiased dynamical trajectories with start and end points located in these predefined stable states. TPS generates this ensemble of transition paths employing a Markov chain Monte Carlo framework Bolhuis et al. (2002). At the core of many path sampling schemes is the so-called shooting move, in which a new path is generated by the integration of the equations of motion (forward and backward in time) starting from a preselected configuration taken from an existing transition path, known as the shooting point. Valid reactive paths that successfully connect the defined states are accepted/rejected in a Metropolis step. Standard TPS has successfully been used on a wide variety of systems and processes Bolhuis and Dellago (2015); Bolhuis and Swenson (2021).
@@
Several methods have been proposed to increase the sampling efficiency by incorporating prior knowledge of the system into the selection of the shooting point using collective variables, such as biased TPS Bolhuis et al. (2002); Juraszek and Bolhuis (2006) or shooting range TPS Jung et al. (2017). Likewise, artificial intelligence-assisted sampling can increase the sampling efficiency significantly Jung et al. (2023); Falkner et al. (2023) and enable efficient calculation of rates Lazzeri et al. (2023). While methods involving prior knowledge about the reaction constitute a powerful branch of TPS algorithm development, there are still many unknown processes where these might not be directly applicable. For such cases, aimless shooting Peters and Trout (2006); Mullen et al. (2015); Peters (2010) and spring shooting Brotzakis and Bolhuis (2016) were introduced as a system-agnostic, yet efficient, alternatives Beckham and Peters (2011); Joswiak et al. (2018); Arjun and Bolhuis (2021); Janoš et al. (2020). While in most other TPS procedures the shooting point is selected freshly from the current path for each trial move, in the aimless and spring shooting methods, the shooting point location (or index) evolves in a certain way from one path to the next, based on knowledge of the shooting point in the previous step. A central concept in these methods is that this allows the shooting algorithm to focus automatically on the barrier region so that the acceptance is improved.
@@
In this study, we revisit several history-based shooting point Monte Carlo methods, including flexible-length aimless shooting and spring shooting. In these methods, we observe that the particular procedure to select shooting points leads to inconsistencies in the detailed balance equations, resulting in an incorrectly sampled path ensemble. This issue arises because the generation of the new shooting point depends on information on the previous shooting, rendering the procedure non-Markovian. Consequently, MCMC theory does not apply to the trajectory space usually considered in the TPS-formalism.
@@
Therefore, our objective is to develop a general formalism that restores Markovianity and considers the evolving shooting point in the detailed balance equation to ensure correct sampling of the desired path distribution. This formalism makes use of an extended state space that includes shooting points in addition to trajectories. We develop the general formalism and apply it to the case of the aimless and spring shooting algorithms. In particular, we show that the performance of the fixed-length methods depends heavily on the total path length. Furthermore, we find a detailed balance-related issue in the original spring shooting method Brotzakis and Bolhuis (2016), which we are able to correct. One of the consequences of treating the algorithms correctly in the extended space formalism is that they are not behaving as efficiently as concluded in the original papers.
@@
The remaining paper is organized as follows: first, we introduce the extended space of paths and shooting indices, which is required to capture the evolution of the index. We then derive detailed balance relations for Monte Carlo moves in this extended space and apply them to regular TPS, aimless shooting and spring shooting. This leads to corrections for flexible-length aimless and spring shooting that are subsequently tested by performing path sampling in a one-dimensional test system.
@@ II.1 Fixed- and Flexible-length Path Ensembles
Assuming two (meta)stable states A and B, we denote the ensemble of paths that connect these states as $P_{\text{AB}}(X)$, where $X=\{x_{1},x_{2},...x_{L}\}$ is a discretized trajectory of length $L$ and each time slice or frame $x$ contains the coordinates (and possibly momenta) of all particles in the system.
The definition of $P_{\text{AB}}(X)$ depends on the desired path ensemble. For instance, a path distribution where all paths are of fixed length $L_{\text{max}}$ can be defined as
$P_{\text{AB}}(X)=\frac{1}{Z_{\text{AB}}}h_{\text{A}}(x_{1})h_{\text{B}}(x_{L})\rho(x_{1})\prod_{i=1}^{L_{\text{max}}-1}p(x_{i}\to x_{i+1})\,,$ (1)
where $h_{\text{A}}(x)$ and $h_{\text{B}}(x)$ are the characteristic functions of the states A and B (unity when $x\in A,B$, zero otherwise), $\rho(x_{1})$ is the stationary distribution generated by the underlying dynamics (e.g. the Boltzmann distribution) and $p(x_{i}\to x_{i+1})$ are the short time transition probabilities. The partition function $Z_{\text{AB}}$ normalizes the distribution.
Another commonly used definition is the flexible-length ensemble, where paths are required to have the start and endpoint in A and B, respectively, while all other points in between should lie outside of A and B:
$P_{\text{AB}}(X)\propto H_{\text{AB}}(X)\prod_{i=1}^{L(X)-1}p(x_{i}\to x_{i+1})\,.$ (2)
Here, $H_{\text{AB}}(X)$ is unity if the path $X$ fulfills all requirements to be considered reactive and zero otherwise, as specified below. Since pathways of different lengths have different measures, one must take special considerations to arrive at a workable distribution. To do so, we consider, in fact, pathways that are of equal length $L_{\text{max}}$ and require that:
@@
If these conditions are met, $H_{\text{AB}}(X)$ is unity and it is zero otherwise. We denote the index of the first point in B as the path length $L(X)$. Note that this way of interpreting the distribution is only formal. In a practical scheme, one halts the integration of the equations of motion when a trajectory enters a stable state. Extending the trajectory (hypothetically) to full length will then result in a valid pathway.
@@ II.2 Extended Space of Paths and Shooting Indices
In path sampling methods like aimless shooting and spring shooting, information on the previous shooting point is used to generate the new shooting point. To take this information into account, we denote with $k$ the index of a (configuration) point $x_{k}$ on the path $X$. This point can either be a shooting point directly or a point that is used to generate the next shooting point depending on the sampling scheme. Independent of the protocol, we refer to $k$ as the shooting index for both cases since, even in the second case, the move is initiated from $k$. Common for both cases is that the index $k$, in some form or the other, evolves from trial to trial. To track the evolution of $X$ and $k$ at the same time in a Monte Carlo procedure and keep the evolution Markovian, we define an extended space consisting of both variables
@@
In order to derive an acceptance probability for the Monte Carlo move, the joint probability density $P(X,k)$ must be specified. Since our goal is the sampling of trajectories distributed according to $P_{\text{AB}}(X)$, we need to require that marginalizing $P(X,k)$ with respect to $k$ should yield $P_{\text{AB}}(X)$:
@@
This condition is naturally fulfilled by defining $P(X,k)$ via the conditional probability $p(k|X)$:
@@
Due to the normalization of the conditional probability, $\sum_{k}p(k|X)=1$, the sum over all $k$ is guaranteed to lead to $P_{\text{AB}}(X)$ as marginal distribution. Apart from this condition, there is total freedom in setting the concrete form of $p(k|X)$, the shooting index distribution.
@@ II.3 Shooting Point Monte Carlo
The goal is now to construct a Markov chain Monte Carlo procedure for sampling the joint probability $P(X,k)=P_{\text{AB}}(X)p(k|X)$ from which we can obtain the desired $P_{\text{AB}}(X)$ by marginalizing with respect to $k$. Imposing detailed balance
@@
to separate the generation probabilities of the new path from the generation probabilities of the new shooting index. For the derivation here, we assume that all pathways are generated using shooting moves with the same underlying dynamics, which conserve the equilibrium distribution. This leads to a cancellation of the path probabilities $P_{\text{AB}}(X)$ and the respective generation probabilities, where we factorize $p_{\text{gen}}$ as in Eq. II.3 Dellago et al. (2002). For the flexible-length path ensemble this leads to:
@@
Below, we will use this equation to derive correct acceptance rules for different TPS-algorithms.
@@ II.4 Shooting Point Densities
To assess the convergence of a path sampling scheme and further understand its inner workings, it is useful to consider the distribution of points on transition paths and the density of points from which shooting moves are initiated. We can derive the density of points on transition paths based on the points $x_{i}$ on paths at indices $i$ and the path density $P_{\text{AB}}(X)$:
@@
where the delta function probes the occurrence of $x$ at position $i$ on a transition path $X$. Note that shorter paths contribute fewer points to this distribution. This density is numerically tractable via a histogram of points or collective variables for complex systems and it is a key indicator for the correct convergence of a path sampling scheme.
Figure 1: Monte Carlo moves in the extended space of path $X$ and shooting index $k$. An index shift (left) generates a new shooting index $k^{\text{o}}\to k^{\text{n}}$ without modifying the current path. The shooting move in the extended space generates a new path $X^{\text{o}}\to X^{\text{n}}$ and, in most cases, a new shooting index $k^{\text{o}}\to k^{\text{n}}$. This is due to the fact that indices may be renumbered after the generation of the new path.
Another distribution that is important to discuss is the distribution of attempted shooting points given that one follows the Monte Carlo scheme described in the previous section. In contrast to $\rho(x|\text{TP})$, this distribution depends on the employed protocol. To obtain a general expression for the distribution of attempted shooting points, we need to consider that $k$ does not need to be the shooting index directly but may also be used to generate the next shooting index. For that, we define $p_{\text{sel}}(\hat{k}|X,k)$ as the probability to initiate shooting from another index $\hat{k}$ given the current index $k$, which can be used to define the (protocol-dependent) shooting point distribution:
@@
For path sampling schemes that do not evolve the shooting point from trial to trial, the new shooting point is often generated independently by directly drawing from $p(k|X)$. In that case, the above expression reduces to
@@
For a uniform shooting index distribution $p(k|X)=\frac{1}{L(X)}$ with independent drawing of shooting indices, as in standard TPS, we arrive at
$\rho(x|\text{SP})=\int\text{d}X\,P_{\text{AB}}(X)\frac{1}{L(X)}\sum_{k=1}^{L(X)}\delta(x-x_{k})\,,$ (15)
which can be transformed to the distribution of points on paths $\rho(x|\text{TP})$ by weighting each path with a factor proportional to $L(X)$.
To gain insights into the relation of the attempted shooting point distributions in equations (II.4), (14) and  (15) to the extended space $(X,k)$, we derive the distribution of points $x_{k}$ located at $k$:
@@
Here, we arrive at an expression equivalent to (14). That means $\rho_{\text{ens}}(x)$, the distribution of points at $k$, is the same as the distribution of attempted shooting points one would obtain if shooting was performed by freshly drawing $k$ each trial from $p(k|X)$. Since $\rho_{\rm ens}(x)$ is independent of the specific shooting scheme yet depends on $p(k|X)$, we use it to assess the convergence of shooting schemes in the extended space $(X,k)$.

### v2

@@ Abstract
Rare event sampling algorithms are essential for understanding processes that occur infrequently on the molecular scale, yet they are important for the long-time dynamics of complex molecular systems. One of these algorithms, transition path sampling (TPS), has become a standard technique to study such rare processes since no prior knowledge on the transition region is required. Most TPS methods generate new trajectories from old trajectories by selecting a point along the old trajectory, modifying its momentum in some way, and then “shooting” a new trajectory by integrating forward and backward in time. In some procedures, the shooting point is selected independently for each trial move, but in others, the shooting point evolves from one path to the next so that successive shooting points are related to each other. To account for this memory effect, we introduce a theoretical framework based on an extended ensemble that includes both paths and shooting indices. We derive appropriate acceptance rules for various path sampling algorithms in this extended formalism, ensuring the correct sampling of the transition path ensemble. Our framework reveals the need for amended acceptance criteria in the flexible-length aimless shooting and spring shooting methods.
@@ I Introduction
Transition path sampling (TPS) is a simulation method that yields unbiased insights into the mechanisms and kinetics of rare events occurring in complex molecular systems. It is applicable to a wide range of processes, including nucleation Menzl et al. (2016), biomolecular reorganization ichi Okazaki et al. (2019); Juraszek and Bolhuis (2006) and chemical reactions Basner and Schwartz (2005); Geissler et al. (2001). Unlike many enhanced sampling techniques, TPS does not require detailed a priori knowledge about the process of interest, other than a definition of the start and end states of the process Bolhuis et al. (2002). TPS samples from a distribution of unbiased dynamical trajectories with start and end points located in these predefined (meta)stable states. This ensemble of transition paths is generated sequentially by employing a Markov chain Monte Carlo framework Bolhuis et al. (2002). At the core of many path sampling schemes is the so-called shooting move, in which a new path is generated by the integration of the equations of motion (forward and backward in time) starting from a preselected configuration taken from an existing transition path, known as the shooting point. Valid reactive paths that connect the start and end states are accepted or rejected in a Metropolis step. Standard TPS has successfully been used on a wide variety of systems and processes Bolhuis and Dellago (2015); Bolhuis and Swenson (2021).
@@
Several methods have been proposed to increase the sampling efficiency of TPS by incorporating prior knowledge of the system into the selection of the shooting point using collective variables, such as biased TPS Bolhuis et al. (2002); Juraszek and Bolhuis (2006) or shooting range TPS Jung et al. (2017). Likewise, artificial intelligence-assisted sampling can increase the sampling efficiency significantly Jung et al. (2023); Falkner et al. (2023) and enable efficient rate calculations Lazzeri et al. (2023). While methods involving prior knowledge about the reaction are useful in many cases, such knowledge is often not available. For such cases, aimless shooting Peters and Trout (2006); Mullen et al. (2015); Peters (2010) and spring shooting Brotzakis and Bolhuis (2016) were introduced as system-agnostic, yet efficient, alternatives Beckham and Peters (2011); Joswiak et al. (2018); Arjun and Bolhuis (2021); Janoš et al. (2020). While in most other TPS procedures the shooting point is selected freshly from the current path for each trial move, in the aimless and spring shooting methods, the new shooting point is chosen based on knowledge of the previous shooting point. As a result, the shooting point evolves in a certain way from one path to the next. A key concept in these methods is that incorporating memory of the shooting point allows the algorithm to focus automatically on the barrier region, thereby increasing the acceptance probability of pathways.
@@
In this study, we revisit several history-based shooting point Monte Carlo methods, including flexible-length aimless shooting and spring shooting. We find that the particular procedures to select shooting points in these methods lead to violations of detailed balance, resulting in an improperly sampled path ensemble. This issue arises because the generation of the new shooting point depends on information on the previous shooting point.
@@
Our objective here is to develop a general formalism that accounts for the evolving shooting point in order to restore detailed balance and ensure correct sampling of the desired path distribution. This formalism makes use of an extended state space that includes shooting points in addition to trajectories. We develop the general formalism and apply it to the case of the aimless and spring shooting algorithms. In particular, we show that the performance of the fixed-length versions of these methods depends strongly on the total path length. Furthermore, we find a detailed balance-related issue in the original spring shooting method Brotzakis and Bolhuis (2016), which we correct in the new formalism. Unfortunately, one of the consequences of treating the algorithms correctly in the extended space formalism is that they are not behaving as efficiently as concluded in the original papers.
@@
The remainder of the paper is organized as follows: first, we introduce the extended space of paths and shooting indices, necessary to capture the evolution of the shooting point. Next, we derive detailed balance relations for Monte Carlo moves in this extended space and apply them to regular TPS, aimless shooting and spring shooting. This formalism leads to corrections of the original flexible-length aimless and spring shooting algorithms, which we then validate by performing path sampling in a one-dimensional test system.
@@ II.1 Fixed- and Flexible-length Path Ensembles
Assuming two (meta)stable states A and B, we denote the ensemble of paths that connect these two states as $P_{\text{AB}}(X)$, where $X=\{x_{1},x_{2},...x_{L}\}$ is a discretized trajectory of length $L$ and each time slice or frame $x$ contains the coordinates (and possibly momenta) of all particles in the system.
The definition of $P_{\text{AB}}(X)$ depends on the desired path ensemble. For instance, a path distribution where all paths are of fixed length $L$ can be defined as
$P_{\text{AB}}(X)=\frac{1}{Z_{\text{AB}}}h_{\text{A}}(x_{1})h_{\text{B}}(x_{L})\rho(x_{1})\prod_{i=1}^{L-1}p(x_{i}\to x_{i+1})\,,$ (1)
where $h_{\text{A}}(x)$ and $h_{\text{B}}(x)$ are the characteristic functions of the states A and B (unity when $x\in\text{A},\text{B}$, zero otherwise), $\rho(x_{1})$ is the stationary distribution generated by the underlying dynamics (e.g., the Boltzmann distribution) and $p(x_{i}\to x_{i+1})$ is the short time transition probability from $x_{i}$ to $x_{i+1}$. The partition function $Z_{\text{AB}}$ normalizes the distribution.
Another commonly used definition is the flexible-length ensemble, in which pathways have different lengths $L(X)$. These paths are required to have the start and endpoint in A and B, respectively, while all other points in between should lie outside of A and B:
$P_{\text{AB}}(X)\propto H_{\text{AB}}(X)\rho(x_{1})\prod_{i=1}^{L(X)-1}p(x_{i}\to x_{i+1})\,.$ (2)
Here, $H_{\text{AB}}(X)$ is unity if the path $X$ fulfills all requirements to be considered reactive and zero otherwise, as specified below.
Since pathways of different lengths have different measures, one must take special precautions to obtain a consistent definition of the transition path ensemble distribution. To do so, we consider, in fact, pathways that are of equal length $L_{\text{max}}$ and require that:
@@
If these conditions are met, $H_{\text{AB}}(X)$ is unity and it is zero otherwise. We denote the index of the first point in B as the path length $L(X)$. Note that this way of interpreting the distribution is only formal. In a practical path sampling scheme, one halts the integration of the equations of motion when a trajectory enters a stable state. Extending the trajectory (hypothetically) to full length $L_{\rm max}$ will then result in a valid pathway.
@@ II.2 Extended Space of Paths and Shooting Indices
In path sampling methods like aimless shooting and spring shooting, information on the previous shooting point is used to generate the new shooting point. To take this information into account, we denote with $k$ the index of a (configuration) point $x_{k}$ on the path $X$. Depending on the sampling scheme, this point can either be a shooting point directly or a point that is used to generate the next shooting point. We refer to $k$ as the shooting index in both cases since, even in the second case, the move is initiated from $k$. Common for both cases is that the index $k$, in some form or the other, evolves from trial to trial.
To track the evolution of $X$ and $k$ at the same time in a Monte Carlo procedure while preserving the Markovianity of the process, we define an extended space that includes both the paths and the shooting indices. Each point $Y$ in this extended space is composed of a path $X$ and a shooting index $k$,
@@
To derive an acceptance probability for the Monte Carlo move, the joint probability density $P(X,k)$ must be specified. Since our goal is to sample the transition path ensemble $P_{\text{AB}}(X)$, we require that marginalizing $P(X,k)$ with respect to $k$ yields $P_{\text{AB}}(X)$,
@@
This condition is naturally fulfilled by defining $P(X,k)$ via the conditional probability $p(k|X)$, which has been described previously in the context of auxiliary variable methods Higdon (1998):
@@
Due to the normalization of the conditional probability, $\sum_{k}p(k|X)=1$, the sum over all $k$ is guaranteed to lead to $P_{\text{AB}}(X)$ as marginal distribution. Apart from this condition, there is total freedom in setting the concrete form of $p(k|X)$, the target distribution for the shooting index $k$.
@@ II.3 Shooting Point Monte Carlo
The goal is now to construct a Markov chain Monte Carlo procedure for sampling the joint probability $P(X,k)=P_{\text{AB}}(X)p(k|X)$. In the subspace of the trajectories $X$ this procedure then samples the desired transition path ensemble $P_{\text{AB}}(X)$. Imposing detailed balance
@@
to separate the generation probabilities of the new path from the generation probabilities of the new shooting index. This factorization implies that first a new path is generated according to $p_{\text{gen}}(X^{\text{n}}|X^{\text{o}},k^{\text{o}})$ based on the old path and the old shooting point. Then, the new shooting point is generated according to $p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})$ based on the old shooting point as well as the old and new path. Another possibility, corresponding to a different factorization, is to first generate the new shooting point and then the new path or further factorize the shooting point generation process.
To further simplify the acceptance probability, we assume that all pathways are generated using shooting moves based on the underlying dynamics, which conserves the equilibrium distribution. Furthermore, we assume that momenta of the shooting point are either unchanged or independently redrawn from the Maxwell-Boltzmann distribution Dellago et al. (2002). This leads to a cancellation of the path probabilities $P_{\text{AB}}(X)$ and the respective generation probabilities, where we factorize $p_{\text{gen}}$ as in Eq. II.3 Dellago et al. (2002). For the flexible-length path ensemble this leads to:
@@
Below, we will use this equation to derive appropriate acceptance rules for different TPS-algorithms.
@@ II.4 Shooting Point Densities
To assess the convergence of a path sampling scheme and further understand its inner workings, it is useful to examine the distribution of points on transition paths, which we note by $\rho(x|\text{TP})$ and the distribution of points from which shooting moves are initiated, indicated by $\rho(x|\text{SP})$ in what follows. For pathways distributed according to $P_{\text{AB}}(X)$, $\rho(x|\text{TP})$ is given by
@@
where the delta function probes the occurrence of $x$ at position $x_{i}$ on a transition path $X$. In the above equation, the notation $\langle\cdots\rangle_{X}$ implies an average over the transition path ensemble $P_{\text{AB}}(X)$. Hence, $\langle L(X)\rangle_{X}$ is the average length of transition pathways. Note that shorter paths contribute fewer points to the distribution $\rho(x|\text{TP})$ defined in Eq. (II.4). For low-dimensional systems, this density can be computed numerically in a TPS simulation by histogramming points on the sampled transition pathways. In higher dimensions, it is more practical to consider the distribution $\rho(r|\text{TP})$ of a collective variable $r(x)$ for points on transition pathways. In either case, the distribution is a key indicator for the correct convergence of a path sampling scheme.
Figure 1: The standard shooting move is irreversible in the extended space framework. Starting from a path $X^{\text{o}}$ and previous shooting point $k^{\text{o}}$ (left), the shooting move is usually divided into two steps: first, a new shooting point $\hat{k}^{\text{o}}$ is picked and, second, the new path $X^{\text{n}}$ is generated from the selected point. On the new path the used shooting point has the index $k^{\text{n}}$. Although the reverse move (right) can recover the old path by selecting the same shooting index $k^{\text{n}}$ again, it is not possible to recover the old shooting index (indicated by the red arrow).
Another distribution that is important to discuss is the distribution of attempted shooting points given that one follows the Monte Carlo scheme described in the previous section. In contrast to $\rho(x|\text{TP})$, which is fully defined by specifying the transition path ensemble $P_{\text{AB}}(X)$, this distribution also depends on the specific protocol employed in the TPS simulation. To obtain a general expression for the distribution of attempted shooting points, we need to take into account that $k$ does not necessarily need to be the shooting index directly, but may also serve to generate an index $\hat{k}$, from which the shooting move is initiated. Denoting with $p_{\text{sel}}(\hat{k}|X,k)$ the probability to select $\hat{k}$ given the current index $k$, the (protocol-dependent) shooting point distribution is given by
@@
Note that the shooting point distribution depends both on the imposed distribution $p(k|X)$ of shooting indices as well as on the specific shooting algorithm as encoded in the selection probability $p_{\text{sel}}(\hat{k}|X,k)$ (and, of course, also on $P_{\text{AB}}(X)$).
For path sampling schemes that do not evolve the shooting point from trial to trial, each new shooting point can be generated independently by directly drawing it from $p(k|X)$, meaning that $p_{\text{sel}}(\hat{k}|X,k)=p(\hat{k}|X)$. In this case, the above expression reduces to
@@
For a uniform shooting index distribution, $p(k|X)=1/L(X)$, with independent drawing of shooting indices, as is done in standard TPS, we obtain the shooting point distribution
$\rho(x|\text{SP})=\int\text{d}X\,P_{\text{AB}}(X)\frac{1}{L(X)}\sum_{k=1}^{L(X)}\delta(x-x_{k})\,.$ (15)
This distribution is very similar to the distribution $\rho(x|\text{TP})$ of points on transition pathways, but it differs from it by a weight factor proportional to $L(X)$. This factor arises because each transition path $X$ contributes $L(X)$ points to the distribution $\rho(x|\text{TP})$, but only one point to the shooting point distributions $\rho(x|\text{SP})$.
To gain further insights into the relation of the attempted shooting point distributions in equations (II.4), (14) and  (15) to the extended space ensemble $P(X,k)$, we determine the distribution of shooting points, not attempted, but as obtained from $P(X,k)$ by integrating/summing over pathways $X$ and shooting indices $k$:
@@
That means that $\rho_{\text{ens}}(x)$, the distribution of points after summing over $k$, is the same as the distribution of attempted shooting points one would obtain if shooting was performed by freshly drawing $k$ each trial from $p(k|X)$. Since $\rho_{\rm ens}(x)$ is independent of the specific shooting scheme yet depends on $p(k|X)$, we use it to assess the convergence of shooting schemes in the extended space $(X,k)$.